40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10962     Entry time: Sat Jan 31 01:34:12 2015
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Not able to engage AO path 

Nothing earth-shattering today.

A few things of note:

  • I checked the coherence (no lock present, just noise) between REFL11_I_IN1 and CM_SLOW_OUT, which are meant to be the same thing when only the REFL1 path is allowed through the CM board.
    • However, at first, there was very little coherence - small band, and only about 0.7 or so.
    • I went inside and jiggled the cables, and also toggled the whitening switches for REFL11 and the CM_SLOW, and after that I had excellent coherence. 
      • I didn't take a coherence spectrum in between those activities, but since the cable connections were all solid, I believe that it may have been a sticky slider -esque problem, and the CM whitening wasn't matched between the analog and digital.
    • I tried two or three times to engage the AO path, but I always lost lock before I was getting any meaningful gain through.
      • I took some TFs remotely with the SR785, but they're totally noise.  I don't even know which sign of the CM board is correct, so no real knkowledge added there, from today.
  • The ~400Hz ringing that we have been seeing, we have been blaming on the PRCL loop.  However, just before my last lockloss I saw gain peaking around 400Hz in the CARM input spectrum. I didn't see if it was also there in the PRCL spectrum.  So, either it is coupling from PRCL to CARM, or CARM itself.  But I think we need to see if we can eek out a little more phase at higher frequencies for both of those loops.  I  just looked, and about 16 seconds before the last lockloss, I see the PRCL loop coupling into the CARM loop.
    • Since yesterday, we have been lowering the PRCL UGF using the servos to be about 70Hz, to give us more gain margin at the high end of the phase bubble, but we still see this ringing. 
  • Two or three times my arm power buildup at 0 CARM offset, 25% MICH offset was at 20 or 21.  Then, a few other times I was only getting about 10 (which is what we have been seeing the last few days.)
    • I'm running the ASS between each lock, although I'm not running it for a full ~2 minutes or so usually. 
    • I think that the reason I was able to get to such high arm powers was excellent alignment, so maybe it's worth sitting and waiting for the ASS to have a full 2 or 3 minutes between locks, even if the ASS error signals look zero-ed out.
    • This is still a factor of 2 lower than we expect for 25% MICH offset, but the whole factor of 5 isn't explained by some mysterious loss.  At least half of it is alignment.
  • The PRCL ASC feedforward still isn't working, but I'd like to try Q's trans qpd ASC soon, with the full lock.  I think the system is ready, but scripts are not, so Q has to be here to run it.

See first plot below for the PRCL->CARM coupling just before lockloss.  The second plot is the corresponding lockloss, where the PRCL loop is starting to see that oscillation again, and it's just barely starting to get into CARM. 

ELOG V3.1.3-